2) A company issues a callable (at par) ten-year, 6% coupon bond with annual coupon payments. The bond can be called at par in one year after release or any time after that on a coupon payment date. On release, it has a price of $104 per $100 of face value. What is the yield to worst of this bond when it is released?

3) A company issues a callable (at par) ten-year, 6% coupon bond with annual coupon payments. The bond can be called at par in one year after release or any time after that on a coupon payment date. On release, it has a yield to maturity of 4.8%, which is below the yield to call. What is the price of this bond per $100 of face value when it is released?

4) A company issues a callable (at par) ten-year coupon bond with annual coupon payments. The bond can be called at par in one year after release or any time after that on a coupon payment date. On release, it has a price of $107 per $100 of face value, and has a yield to call of 3.5%. What is the bond's coupon rate?

5) A bond has a face value of $10,000 and a conversion price of $37.74. The stock is currently trading at $38.80. What is the conversion ratio?

7) A firm requires an investment of $20,000. The firm's debt cost of capital is 5%, and its return on equity is 12%. If the firm's pre-tax WACC is 7.8%, how much equity did the firm use for its investment?
Answer #1

Solution:

Q2.

It is given that the

the par value = $100.

Years to maturity = 10 years.

Coupon rate = 6%. Annual coupon payment = 6% *100 = $6  

Present value of bond = $104 and Future value is = $100

This bond is callable after one year hence year remains till bond is callable = 1.

We can use two methods to calculate the answer.

Method 1:

Yield to call = {Coupon + ( Face Value - Present Value or call price) / Years till call } /{ (face value + Present Value or call price)/2}

Yield to call = {6 + ( 100 - 104) / 1 } /{ (100 + 104)/2} = 2/102 = 1.96%

This formula gives approximate answer

Method 2: We can use excel to find the exact rate by using rate formula

Yield to worst = RATE (period, coupon, present value, future value) = RATE(1,6,-104,100) = 1.92%

1.92%]=RATE(1,6,-104,100)

Correct answer is 1.92%